OpenStudy (anonymous):

what is the formula to find the acceleration ? is it final velocity (v) - initial velocity (u) divided by time taken ?

OpenStudy (anonymous):

Acceleration There are a few. The most famous is a = F/m, where F is the net force applied to a mass, m. Acceleration is also the change in velocity, (Delta-V), divided by the change in time, (Delta-t). So, a = Δv/Δt. For example, if an object's velocity changes from 10 meters per second to 20 meters per second in five seconds, its acceleration is (20-10)/5 = 2 meters per second.